### v1

@@
11institutetext: Department of Mechanical Engineering, Texas A&\&&M University
11email: {kasiv, ssaripalli}@tamu.edu
22institutetext: DEVCOM Army Research laboratory, Adelphi, MD, USA
22email: {felix.a.sanchez12.ctr, timothy.j.overbye.ctr, jason.m.gregory1.civ}@army.mil
Trailblazer: Learning offroad costmaps for long range planning
Kasi Viswanath 11    Felix Sanchez 22    Timothy Overbye 22    Jason M. Gregory 22    Srikanth Saripalli 11
@@ 1 Introduction
Recent advancements have introduced imitation learning [3][4] and Vision Language Models [5] to optimize cost-maps through human demonstrations, reducing manual intervention. However, many of these methods still depend on pre-trained convolutional neural networks (CNNs) trained on manually crafted cost-maps, further refined using reinforcement learning algorithms.
Trailblazer offers a novel alternative by directly converting multi-modal sensor data into cost-maps. Instead of relying on pre-trained models or manual annotations, Trailblazer learns cost-map generation from expert driven demonstrations. This approach simplifies the process, eliminates manual tuning, and enhances the system’s adaptability and efficiency in real-world scenarios.
@@ 2 Methodology
Trailblazer methodology is divided into two key stages: (i) data collection and pre-processing, where satellite imagery and LiDAR data are gathered and prepared; (ii) Trailblazer framework and training, which involves the application of a sophisticated architecture for feature extraction and path planning.
@@ 2.1 Data collection and pre-processing
The Trailblazer architecture utilizes satellite/aerial imagery and airborne LiDAR point cloud data as primary inputs. This study sources data predominantly from the United States Geological Survey (USGS) and the Environmental Systems Research Institute (ESRI) via ArcGIS servers [6]. ESRI provides high-resolution satellite imagery with a spatial resolution of up to 30 cm per pixel, while the USGS LiDAR Explorer delivers point cloud data with a density exceeding 8 points per square meter. Imagery data contributes essential information about land cover, terrain, transportation networks, and water bodies, while LiDAR data offers detailed insights into terrain features such as slope profiles, height measurements, and intensity maps.
@@
Figure 1: Inputs to Trailblazer from Texas A&M RELLIS test site.
@@ 2.2.1 Encoder-Decoder:
Figure 2: Trailblazer framework architecture. The whole network is differentiable.
@@ 3.2 Trailblazer
Figure 3: Semantic ground truths and predictions from Segformer.
Figure 4: Training and validation loss curves of Trailblazer.
To enable sim-to-real transfer, we retrained the simulation-trained Trailblazer model using real-world trajectory data. The model was evaluated on various real-world overhead datasets obtained from the USGS, with an example illustrated in Figure 5.
3.2.1 Planned experiments:
OpenStreetMap (OSM) [14] provides GPS trajectories of vehicles driven across various locations, which we leverage to further train Trailblazer for enhanced generalization across diverse geographic regions. These trajectories, exemplified in Figure 2, serve as valuable data to improve the model’s adaptability. Concurrently, real-time navigation experiments are being conducted, where Trailblazer functions as a global planner in conjunction with a local planner. This integrated framework aims to enhance navigation efficiency and reliability, ensuring seamless movement between points in diverse environments.
@@ 4 Experimental Insights
We experimented with replacing 3D LiDAR scans with Digital Elevation Maps (DEMs), which are widely available. While DEMs produced costmaps of comparable quality to 3D LiDAR, their effectiveness depended heavily on resolution. DEMs with resolutions below 1/3 arc-second resulted in subpar costmaps and slope distortions.
Trailblazer, initially a local planning algorithm for off-road terrains, now supports both local and global planning by dynamically updating the global costmap with real-time onboard data. For instance, if overhead data indicates a dry, traversable riverbed but local inspection reveals otherwise, the global costmap adapts, enabling a new optimal trajectory and ensuring robust planning in dynamic environments.

### v2

@@
Abstract
Autonomous navigation in off-road environments remains a significant challenge in field robotics, particularly for Unmanned Ground Vehicles (UGVs) tasked with search and rescue, exploration, and surveillance. Effective long-range planning relies on the integration of onboard perception systems with prior environmental knowledge, such as satellite imagery and LiDAR data. This work introduces Trailblazer, a novel framework that automates the conversion of multi-modal sensor data into costmaps, enabling efficient path planning without manual tuning. Unlike traditional approaches, Trailblazer leverages imitation learning and a differentiable A* planner to learn costmaps directly from expert demonstrations, enhancing adaptability across diverse terrains. The proposed methodology was validated through extensive real-world testing, achieving robust performance in dynamic and complex environments, demonstrating Trailblazer’s potential for scalable, efficient autonomous navigation. https://github.com/unmannedlab/Trailblazer
keywords:
Off-Road, Long-Range planning, Navigation, Autonomous Driving
@@ 1 Introduction
Previous work, [3] proposed an approach to estimate a cost function for path planning using human demonstrations. This algorithm makes use of overhead data sources like satellite images and aerial LiDAR scans, processing them directly without going through the usual feature extraction steps. To manage the complexities of different scenes, linear regressors were applied. However, the absence of feature extraction can lead to a loss of important information. This may hinder the algorithm’s effectiveness, especially in more complicated environments.
Recent advancements have introduced deep reinforcement learning [4][5] and Vision Language Models [6] to optimize costmaps through human demonstrations, reducing manual intervention. However, many of these methods still depend on pre-trained convolutional neural networks (CNNs) trained on manually crafted costmaps, further refined using reinforcement learning algorithms.
Trailblazer offers a novel alternative by directly converting multi-modal sensor data into costmaps. Instead of relying on pre-trained models or manual annotations, Trailblazer learns cost-map generation from expert driven demonstrations. This approach simplifies the process, eliminates manual tuning, and enhances the system’s adaptability and efficiency in real-world scenarios for global planning.
@@ 2 Methodology
Trailblazer methodology consists of two stages: (i) data collection and pre-processing of satellite imagery and LiDAR data; (ii) Trailblazer framework and training for feature extraction and path planning.
@@ 2.1 Data collection and pre-processing
The Trailblazer architecture utilizes satellite/aerial imagery and airborne LiDAR point cloud data as primary inputs. The Trailblazer architecture makes use of satellite and aerial images, along with airborne LiDAR point cloud data, as its main sources of information. For this study, the data comes mainly from two reliable sources: the United States Geological Survey (USGS)[7] and the Environmental Systems Research Institute (ESRI), specifically through ArcGIS servers [8]. ESRI provides high-quality satellite images that can achieve a spatial resolution of up to 30 centimeters per pixel. On the other hand, the USGS LiDAR Explorer offers point cloud data with a density that goes beyond 8 points for every square meter.
Figure 1: Inputs to Trailblazer from Texas A&M RELLIS test site. The satellite image is of 30 $cmperpixel$ resolution used to generate semantic map from SegFormer. Overhead LiDAR data is used to generate the height, slope and intensity maps. The RELLIS test site covers an area of $1.5km*1.5km$.
@@
Semantic projection is derived using SegFormer[9] on satellite imagery, classifying pixels by an ascending risk factor, from traversable to water. Segmented masks are projected onto LiDAR point clouds and mapped onto a grid. For each grid cell, the highest class index is assigned as its value. Geometric characteristics are extracted from LiDAR data. An average height map is created by calculating the average height of all points in each grid cell. LiDAR intensity is also analyzed on a per-cell basis. The average intensity per grid cell helps distinguish terrain types (e.g., trees, vegetation, barren land) based on surface characteristics. LiDAR intensity, also averaged per cell, helps differentiate various terrain types (e.g., trees, vegetation, barren land) based on surface characteristics.
We also work with Digital Elevation Maps(DEMs) for areas where LiDAR data is not available. The DEMs are transformed into a 2.5D point cloud, which is based on the resolution of the scan. After this conversion, we process the data similarly to how we handle LiDAR data, to create the necessary inputs for Trailblazer.
@@ 2.2.1 Encoder-Decoder:
Figure 2: Trailblazer framework architecture. The backbone consists of feature fusion and feature extraction branch generating the costmap. The Neural A* planner generates the path from the costmap compared against ground truth to compute loss.
@@ 3.1 Semantic Segmentation
Figure 3: Semantic ground truths and predictions from Segformer.
Figure 4: Training and validation loss curves of Trailblazer.
@@ 3.2 Trailblazer
To enable sim-to-real transfer, we retrained the simulation-trained Trailblazer model using real-world trajectory data. OpenStreetMap (OSM) [16] provides GPS trajectories of vehicles driven across various locations, which we leverage to further train Trailblazer for enhanced generalization across diverse geographic regions. These trajectories, exemplified in Figure 5, serve as valuable data to improve the model’s adaptability.The model was evaluated on various real-world overhead datasets obtained from the USGS, with an example illustrated in Figure 5.
3.2.1 Real-world Experiments:
Real-time navigation experiments were conducted at sites in North Carolina in association with DARPA RACER program, where Trailblazer functions as a global planner in conjunction with a local planner.
@@ 3.2.1 Real-world Experiments:
Figure 6: LiDAR vs DEM: (a) Satellite image from a test site (b) Costmap generated from LiDAR data with 4 $points/m^{2}$ resolution., (c) Costmap generated from DEM with a $1/3arc-sec$ resolution.
@@ 4 Experimental Insights
Trailblazer also includes capbility to generate costmaps from Digital Elevation Maps(DEMs) replacing 3D LiDAR data. While DEMs produced costmaps of comparable quality to 3D LiDAR, their effectiveness depended heavily on resolution. DEMs with resolutions below 1/9 arc-second resulted in subpar costmaps and slope distortions. Figure 6 compares the quality of costmaps generated from LiDAR and DEMs. Since LiDAR data isn’t globally available, we aimed to utilize widely mapped DEMs.
While handling data from different sources, one must be careful when transforming coordinates between various datums. An incorrect transformation can lead to misalignment. This misalignment can result in inaccurate projections and costmaps that do not line up properly.
Trailblazer started as a local planning algorithm[17] designed for navigating off-road terrains. Over time, it has evolved to handle not just local planning, but also global planning. This advancement is made possible by the algorithm’s ability to dynamically update the global costmap with real-time data collected onboard. For instance, consider a situation where overhead data may indicate the presence of a dry riverbed as a viable path. However, a local inspection might show that this specific area is, in fact, not traversable. In such cases, the global costmap can adjust itself based on the latest information, which enables Trailblazer to determine the best trajectory possible. This feature fosters strong and effective planning in environments that are constantly changing.
5 Conclusion
In this paper, we presented Trailblazer, a novel imitation learning framework designed for long-range planning in off-road terrains. By leveraging multi-modal sensor data, including satellite imagery and LiDAR, Trailblazer generates costmaps for efficient path planning without the need for manual tuning. Extensive real-world testing validates Trailblazer’s robust performance, demonstrating its potential for scalable and efficient autonomous navigation. The framework also showcases the utility of open-source data from sources like USGS, ESRI, and OpenStreetMap, proving that such data is sufficient for robust urban and off-road navigation.
Trailblazer’s ability to generate costmaps from Digital Elevation Maps (DEMs) offers a viable alternative when LiDAR data is unavailable, although the effectiveness depends on the resolution of the DEMs. Looking ahead, we plan to adjust our framework so that it can utilize high-resolution satellite images alongside low-resolution Digital Elevation Maps. Our goal is to create costmaps that are similar in quality to those made with LiDAR data. This approach will allow us to produce costmaps using data that is more widely accessible all over the world.
6 Acknowledgments
Research was sponsored by DEVCOM Army Research Laboratory (ARL) and was accomplished under Cooperative Agreement Number W911NF-21-2-0064. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the ARL or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.